HUBBLE SPACE TELESCOPE DAILY REPORT #2534 PERIOD COVERED: 0000Z (UTC) 01/06/00 - 0000Z (UTC) 01/07/00 Daily Status Report as of 007/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Six Sets of FGS 8532 (White Dwarf and M-Stars in the Solar Neighborhood; Revealing Binarity) FGS-1 was used to conduct a high angular resolution survey with FGS1r to detect binarity among nearby M dwarfs and white dwarfs. As reported in yesterday's report and HSTAR 7465, while the acquisition for the first iteration of this proposal succeeded, no astrometry was performed. Otherwise, the observations completed with no further reported problems. 1.2 Completed STIS/MA1 8510 (FUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA1) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.3 Completed Two Sets of STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.4 Completed STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.5 Completed Two Sets of STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.6 Completed WF/PC-2 8492 (UV Monitor for -88C Cooldown) The WF/PC-2 was used to conduct the UV monitor after the first cooldown to -88C. Throughput measurements are also made using UV filters and UV crossed with long-pass filters to determine red leak contribution. -These are used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror. The proposal completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 11 Successful: 11 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 29 Successful: 29 2.3 Operations Notes: The Servicing Mission Observatory Verification (SMOV) program is continuing. Using ROP SR-6 as a guide, there were four SSR-1 memory location dumps. The SSR pointer was set twice as directed by ROP SR-3. The SSA transmitter was turned on and off at 006/1453Z per ROP IC-2. SSAT-2 was power cycled at 006/2046Z. The 486 was dumped, and the results were compared successfully, at 006/2136Z. ROP DF-06A was used. Using ROP NS-3, the NSSC-1 status buffer was dumped and reset at 007/0105Z. The Vehicle Disturbance Test (VDT) procedure was again initiated at 007/0940Z. The results have been nominal. The next iteration of this test will be conducted tomorrow. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW